OpenStudy (anonymous):

solve the equatkion 7/x+4-5/x-4=12/-x^2-16

OpenStudy (anonymous):

\[\frac{7}{x+4}-\frac{5}{x-4}=\frac{12}{-x^2-16}\]

OpenStudy (anonymous):

are you sure it is not \[\frac{7}{x+4}-\frac{5}{x-4}=\frac{12}{x^2-16}\]

OpenStudy (anonymous):

the bottom one is right